# Cosmology in Modified $f(\mathcal{G})$ Gravity: A Late Time Cosmic Phenomena

### v1

@@ IV Stability analysis of dynamical systems
Critical Point X Y Z V Existence $\omega_{\text{eff}}$ $q$
$\mathcal{P}_{1}=(x_{1},y_{1},z_{1},v_{1})$ $x_{1}$ $4x_{1}$ $-2$ $1+24x_{1}$
$x_{1}\neq 0$,   $m=\frac{1}{4}$
$\frac{1}{3}$ $1$
$x_{2}\neq 0$,   $m=\frac{1}{4}$
00 $\frac{1}{2}$
$\mathcal{P}_{3}=(x_{3},y_{3},z_{3},v_{3})$ $x_{3}$ $3x_{3}$ $-\frac{3}{2}$ 00
$x_{3}\neq 0$,   $m=\frac{1}{4}$
$-1+\frac{-1-8x_{3}}{8x_{3}}$ $-1+\frac{-1-8x_{3}}{12x_{3}}$
$\mathcal{P}_{4}=(x_{4},y_{4},z_{4},v_{4})$ $\frac{-1}{16}$ $\frac{-1}{8}$ $-1$ 00
$2m^{2}-m\neq 0$
$-\frac{1}{3}$ 00
$\mathcal{P}_{5}=(x_{5},y_{5},z_{5},v_{5})$ $\frac{m}{-4+8m}$ 00 00 00
$(-1+2m)(-1+4m)\neq 0$, $m\neq 0$
$-1$ $-1$
$\mathcal{P}_{1}$ 00 $1+24x_{1}$ $-24x_{1}$ Never
$\mathcal{P}_{2}$ $1+20x_{2}$ 00 $-20x_{2}$ Never
$\mathcal{P}_{3}$ 00 00 $1$ $x_{3}<-\frac{1}{16}$vee $x_{3}>0$
$\mathcal{P}_{4}$ 00 00 $1$ Never
$\mathcal{P}_{5}$ 00 00 $1$ Always
Table 4: The density parameters associated with the critical point.
Critical Point Eigenvalues Stability
$\mathcal{P}_{3}$$\left\{0,-6-\frac{1}{4x_{3}},-5-\frac{1}{4x_{3}},-4-\frac{1}{8x_{3}}\right\}$ Stable for $x_{3}<-\frac{1}{20}$vee $x_{3}>0$
$\mathcal{P}_{4}$$\left\{-2,-2,-1,4+\frac{2}{-1+2m}\right\}$ Stable for $\frac{1}{4}<m<\frac{1}{2}$
$\mathcal{P}_{5}$$\left\{-4,-3,\frac{3m-6m^{2}-\sqrt{(1-2m)^{2}(25m-4)m}}{2m(2m-1)},\frac{3m-6m^{2}+\sqrt{(1-2m)^{2}(25m-4)m}}{2m(2m-1)}\right\}$ Stable for $\frac{4}{25}\leq m<\frac{1}{4}$
Table 5: Eigenvalues and stability regime.
Figure 7: 2D phase portrait for the dynamical system.
The 2D phase portraits presented in Fig. 7 illustrate the behavior of the dynamical system under the condition $m=\frac{1}{4}$ by projecting the trajectories in the $XY$ and $XV$ planes. These visualizations provide insights into the stability and nature of the critical points $\mathcal{P}_{1}$ to $\mathcal{P}_{5}$.
@@ 00 $\frac{1}{2}$
  • Critical Point $\mathcal{P}_{3}$: The critical point $\mathcal{P}_{3}$ is shown on the $XV$ plane in Fig. 7. The deceleration parameter’s value at this critical point, along with the equation of state parameters, relies on the coordinates $x_{3}$. The value of $q=-1+\frac{-1-8x_{3}}{12x_{3}}$, and $\omega_{\text{eff}}=-1+\frac{-1-8x_{3}}{8x_{3}}$. This crucial point signifies a period dominated by standard cold dark matter with $\Omega_{\text{DE}}=1$, which can account for the current cosmic acceleration when ($x_{3}<\frac{-1}{16}\vee x_{3}>0$). The trajectories suggest the presence of a saddle point, similar to $\mathcal{P}_{2}$. The system is attracted along specific directions (stable manifolds) and repelled in others (unstable manifolds). This behavior aligns with the eigenvalue analysis, indicating the conditional stability of $\mathcal{P}_{3}$.
  • Critical Point $\mathcal{P}_{4}$: The critical point $\mathcal{P}_{4}$ is illustrated on the $XY$ plane in Fig. 7. It is important to note that the critical point $\mathcal{P}_{4}$, where $q=0$ and $\omega=\frac{-1}{3}$, does not depict accelerating expansion. This critical point occurs at $2m^{2}-m\neq 0$ and is stable when $\frac{1}{4}<m<\frac{1}{2}$. The deceleration parameter solely relies on the variable $Z=\frac{\dot{H}}{H^{2}}$. In this context, when $Z=-1$, it signifies that the contribution from the GB invariant $\mathcal{G}=24H^{4}(Z+1)$ term vanishes, which could explain why it doesn’t demonstrate the transition phase, as depicted in Fig. 7. The phase portrait shows trajectories spiralling inwards towards $\mathcal{P}_{4}$, indicating that it is a stable spiral. This suggests that small perturbations will cause the system to oscillate while eventually converging to $\mathcal{P}_{4}$. The eigenvalues, consisting of negative real parts, confirm the stable nature of this critical point.
  • Critical Point $\mathcal{P}_{5}$: The critical point $\mathcal{P}_{5}$ is depicted on the $XY$ plane in Fig. 7. The critical point $\mathcal{P}_{5}$ represents the de-Sitter solutions with $\Omega_{\text{DE}}=1$, $\Omega_{\text{m}}=0$, $\Omega_{\text{r}}=0$, indicating the current accelerated expansion of the Universe. This de-sitter solution is valid at the critical point $\mathcal{P}_{5}$ within the parameter range $(-1+2m)(-1+4m)\neq 0$ and $m\neq 0$. Consequently, the value of $\omega_{\text{eff}}=q=-1$ highlights the significance of this critical point in describing the current dynamics of the Universe. The trajectories show a clear spiral structure converging towards $\mathcal{P}_{5}$, indicating a stable focus. This implies that the system will exhibit damped oscillations as it approaches $\mathcal{P}_{5}$. The stability analysis of $\mathcal{P}_{5}$ supports this observation, showing that the real parts of the eigenvalues are negative, thus confirming stability.
@@
Figure 8: Evolution of the relative energy density of dark matter ($\Omega_{\text{m}}$), radiation ($\Omega_{\text{r}}$) and DE ($\Omega_{\text{DE}})$.The vertical dashed red line indicates the present cosmological time.
In the system described by equations (33), it is possible to perform numerical integration with suitable initial conditions to capture the complete cosmological evolution across different epochs. Current cosmological parameter measurements [59] suggest that the Universe is almost flat. For this specific example, we use the initial conditions $X=10^{1.45}$, $Y=-1\times 10^{-5}$, $Z=-0.1$, and $V=10^{-4}$. The behavior observed aligns with current cosmic observations regarding the evolution of density parameters. The resulting current densities are $\Omega_{\text{m}}(0)=0.3$, $\Omega_{\text{DE}}(0)=0.69$, and $\Omega_{\text{r}}(0)=10^{-4}$. By integrating equation (30) using the summarized initial conditions, we obtain numerical solutions for the density parameters $\Omega_{\text{m}}$, $\Omega_{\text{r}}$, and $\Omega_{\text{DE}}$, as shown in Fig. 8. These results reveal that the Universe evolves through a radiation-dominated phase at early times $(q=1)$. Subsequently, it transitioned into a matter-dominated phase with a deceleration parameter $\frac{1}{2}$. Currently, it is moving into an exponentially accelerating epoch with a deceleration parameter of $-1$.
V Conclusion
@@
We investigated the cosmological properties of the $f(\mathcal{G})$ model in the presence of matter fields. We assumed non-relativistic pressureless matter and neglected the late-time contribution of the radiation fluid. By numerically solving the first Friedmann equation, we determined the redshift behavior of the Hubble parameter. We used the $\Lambda$CDM model to establish appropriate initial conditions for $H(z)$ and its derivatives. Subsequently, we utilized the most recent low-redshift observations to compare our theory directly with the model-independent predictions of the cosmic expansion. Specifically, we employed a Bayesian analysis with the MCMC method, incorporating Pantheon+ and CC datasets. By assuming uniform prior distributions, we obtained constraints on the free parameters of the model at the 1-$\sigma$ and 2-$\sigma$ confidence levels. This enabled us to reconstruct the cosmological evolution of the Hubble expansion rate and the total effective equation of state parameter. Our analysis indicates that the $f(\mathcal{G})$ model effectively accounts for the current acceleration of the Universe without the need for $\Lambda$. However, upon closer analysis and comparison with the predictions of the standard cosmological scenario, it becomes evident that the $f(\mathcal{G})$ model exhibits significant deviations from $\Lambda$CDM as the redshift increases, demonstrating its inability to describe a standard matter-dominated era.
@@
In the second phase of our study, we conducted a dynamical system analysis, focusing on the type of $f(\mathcal{G})$ function under consideration. This analysis allowed us to study the global behavior and stability of the cosmological model. We made some interesting preliminary findings about the finite phase space of a power-law class of the GB gravity model. Equations (33) describe the dynamical system for the power law $f(\mathcal{G})$ gravity model. In Table 3, we provided critical points, existing conditions, effective equation of state (EoS), and deceleration parameters for the autonomous system. Table 4 presents density parameter values for the acceleration phase. We obtained a total of five critical points, with four of them ($\mathcal{P}_{2}$, $\mathcal{P}_{3}$, $\mathcal{P}_{4}$, $\mathcal{P}_{5}$) being stable and the remaining one ($\mathcal{P}_{1}$) being unstable.

### v2

@@ 00 $\frac{1}{2}$
Figure 8: Two-dimensional phase portrait for the dynamical system.
The two-dimensional phase portraits shown in Fig. 8 depict the dynamics of the system for $m=\frac{1}{4},0.3,0.18$ by mapping the trajectories onto the $XY$ and $XV$ planes. These visualizations provide insights into the stability and nature of the critical points $\mathcal{P}_{1}$ to $\mathcal{P}_{4}$.
@@
  • Critical Point $\mathcal{P}_{3}$: The critical point $\mathcal{P}_{3}$ is illustrated on the $XY$ plane in Fig. 8. It is important to note that the critical point $\mathcal{P}_{3}$, where $q=0$ and $\omega=\frac{-1}{3}$, does not depict accelerating expansion. This critical point occurs at $2m^{2}-m\neq 0$ and is stable when $\frac{1}{4}<m<\frac{1}{2}$. The deceleration parameter solely relies on the variable $Z=\frac{\dot{H}}{H^{2}}$. In this context, when $Z=-1$, it signifies that the contribution from the GB invariant $\mathcal{G}=24H^{4}(Z+1)$ term vanishes, which could explain why it doesn’t demonstrate the transition phase, as depicted in Fig. 8. The phase portrait shows trajectories spiralling inwards towards $\mathcal{P}_{3}$, indicating that it is a stable spiral. This suggests that small perturbations will cause the system to oscillate while eventually converging to $\mathcal{P}_{3}$. The eigenvalues, consisting of negative real parts, confirm the stable nature of this critical point.
  • Critical Point $\mathcal{P}_{4}$: The critical point $\mathcal{P}_{4}$ is depicted on the $XY$ plane in Fig. 8. The critical point $\mathcal{P}_{4}$ represents the de-Sitter solutions with $\Omega_{\text{DE}}=1$, $\Omega_{\text{m}}=0$, $\Omega_{\text{r}}=0$, indicating the current accelerated expansion of the Universe. This de-sitter solution is valid at the critical point $\mathcal{P}_{4}$ within the parameter range $(-1+2m)(-1+4m)\neq 0$ and $m\neq 0$. Consequently, the value of $\omega_{\text{eff}}=q=-1$ highlights the significance of this critical point in describing the current dynamics of the Universe. The trajectories show a clear spiral structure converging towards $\mathcal{P}_{4}$, indicating a stable focus. This implies that the system will exhibit damped oscillations as it approaches $\mathcal{P}_{4}$. The stability analysis of $\mathcal{P}_{4}$ supports this observation, showing that the real parts of the eigenvalues are negative, thus confirming stability.
@@
Figure 9: Evolution of the relative energy densities of dark matter $\Omega_{\text{m}}$, radiation $\Omega_{\text{r}}$, and DE $\Omega_{\text{DE}}$. The thick line represents the evolution of the density parameter for the $f(\mathcal{G})$ model, while the dashed line shows the evolution for the $\Lambda$CDM model.
In the system described by equations (37), it is possible to perform numerical integration with suitable initial conditions to capture the complete cosmological evolution across different epochs. Current measurements of cosmological parameters (Aghanim et al., 2020) suggest that the Universe is almost flat. For this specific example, we use the initial conditions $X=10^{14}$, $Y=-1.2\times 10^{14}$, $Z=0.005$, and $V=8.2\times 10^{-5}$ and model parameter $m=0.18$. The behavior observed aligns with current cosmic observations regarding the evolution of density parameters. By integrating equation (34) using the summarized initial conditions, we obtain numerical solutions for the density parameters $\Omega_{\text{m}}$, $\Omega_{\text{r}}$, and $\Omega_{\text{DE}}$, as shown in Fig. 9. These results reveal that the Universe evolves through a radiation-dominated phase at early times $(q=1)$. Subsequently, it transitioned into a matter-dominated phase with a deceleration parameter $\frac{1}{2}$. Currently, it is moving into an exponentially accelerating epoch with a deceleration parameter of $-1$. The $f(\mathcal{G})$ model represents a cosmological scenario in which the Universe undergoes successive eras of radiation domination, matter domination, and currently, DE domination. Our results indicate that the point where matter and radiation contribute equally is slightly higher than in the $\Lambda$CDM model. The behavior of the model is consistent with current cosmic observations on the evolution of density parameters. The current densities are approximately $\Omega_{\text{m}}\approx 0.3$, $\Omega_{\text{DE}}\approx 0.7$, and $\Omega_{\text{r}}\approx 10^{-4}$. Similar behavior in the evolution of density parameters have been noted in the literature (Granda (2020)).
5 Conclusion
@@
We investigated the cosmological properties of the $f(\mathcal{G})$ model in the presence of matter fields. We assumed non-relativistic pressureless matter and neglected the late-time contribution of the radiation fluid. By numerically solving the first Friedmann equation, we determined the redshift behavior of the Hubble parameter. We used the $\Lambda$CDM model to establish appropriate initial conditions for $H(z)$ and its derivatives. Subsequently, we utilized the most recent low-redshift observations to compare our theory directly with the model-independent predictions of the cosmic expansion. Specifically, we used a Bayesian analysis with the MCMC method, analysing using the Pantheon+ and CC datasets separately. By assuming uniform prior distributions, we obtained constraints on the free parameters of the model at the 1-$\sigma$ and 2-$\sigma$ confidence levels. This enabled us to reconstruct the cosmological evolution of the Hubble expansion rate and the total effective EoS parameter. Our analysis indicates that the $f(\mathcal{G})$ model effectively accounts for the current acceleration of the Universe without the need for $\Lambda$. However, upon closer analysis and comparison with the predictions of the standard cosmological scenario, it becomes evident that the $f(\mathcal{G})$ model exhibits significant deviations from $\Lambda$CDM as the redshift increases, demonstrating its inability to describe a standard matter-dominated era. Our analysis identifies a notable discrepancy between the $H_{0}$ values obtained from the CC sample and the Pantheon datasets, as seen in Fig. 3, highlighting the ongoing $H_{0}$ tension in cosmology. This result emphasizes the necessity for further research into possible systematic errors or new physics to resolve this issue. Addressing these discrepancies is essential for enhancing our understanding of the expansion rate of the Universe.
@@
In the second phase of our study, we conducted a dynamical system analysis, focusing on the type of $f(\mathcal{G})$ function under consideration. This analysis has enabled us to assess the global behavior and stability of the cosmological model. It provided insights into the critical points associated with the model and their characteristics, which could be relevant to observable cosmology and the evolution of the Universe. Our findings revealed stable critical points describing the late-time cosmic accelerated phase. This indicates non-standard matter and radiation-dominated eras of the Universe. Interestingly, our results align with the standard quintessence model on $z>0$. We made notable preliminary discoveries regarding the finite phase space of a power-law class of the GB gravity model. In addition, the equations describing the dynamical system for the power law $f(\mathcal{G})$ gravity model are provided in Equations (37). Furthermore, Table 3 includes critical points, existing conditions, effective EoS, and deceleration parameters for the autonomous system. In contrast, Table 4 presents density parameter values for the acceleration phase. In total, we identified four critical points, three being stable ($\mathcal{P}_{2}$, $\mathcal{P}_{3}$, $\mathcal{P}_{4}$) and one unstable ($\mathcal{P}_{1}$). Finally, Fig. 9 shows the proficiency of the model in depicting the evolution of dark matter, radiation, and DE densities, effectively capturing the transitions through various cosmic epochs and reinforcing the model’s robustness in explaining late-time cosmic phenomena.
Data Availability
There are no associated data with this article. No new data were generated or analysed in support of this research.
Acknowledgments
S V L would like to express gratitude for the financial support provided by the University Grants Commission (UGC) through the Senior Research Fellowship (UGC Reference No. 191620116597) to carry out the research work. BM acknowledges the support of the Council of Scientific and Industrial Research (CSIR) for the project grant (No. 03/1493/23/EMR II). We sincerely appreciate the esteemed referee and editor for their insightful suggestions, which have greatly enhanced the quality and presentation of our research.
The authors declare that there is no conflicts of interest.